# Automatic dataset shift identification to support root cause analysis of AI performance drift

### v2

@@ Results
To perform shift identification, we leverage the fact that output-based and feature-based shift measures detect different types of shifts to precisely identify the shift present in the test dataset, following a decision logic detailed in Methods Dataset shift identification and illustrated in Fig. 2.
@@
In-depth evaluation results in Figs. 5 and 6 demonstrate that our novel shift identification framework is capable of distinguishing between prevalence shifts, covariate shifts and mixed shifts with high accuracy across all datasets and types of shifts. Overall, more subtle shifts are best detected with larger test sets whereas larger shifts can be detected with smaller test sets. For prevalence shifts, the average shift identification rate, across datasets and shift levels, is $85\%$ with 500 test images and reaches $95\%$ with 1,000 test images (Fig. 5 a-d). Similarly, for covariate shifts caused by acquisition shifts, when using a test set size of 500 images (and 250 test exams on EMBED), the identification accuracy is greater than $80\%$ for any shift level and dataset, with an average identification accuracy of $89\%$ across datasets and shift levels (Fig. 5 g-i). For the RETINA dataset, identification rates already reach $100\%$ with a test set as small as 250 images. When the covariate shift is induced by gender shifts, covariate shift is detected with identification accuracies greater than $90\%$ for both PadChest and RSNA Pneumonia, with a test set size of 500 images, for all but one shift level (Fig. 5 e-f). Results in Fig. 6, show that the framework is also able to accurately distinguish between cases of covariate shift only and cases of covariate and prevalence shifts. For these mixed shifts, shifts are identified with increasingly high accuracy as the test set size increases. With a test set of $1,000$ images, mixed gender and prevalence shifts are correctly identified as mixed shifts with an average accuracy of $97\%$ for RSNA Pneumonia and $75\%$ for PadChest. Mixed shifts induced by acquisition and prevalence shifts are detected with an average $100\%$ accuracy for the RETINA dataset, $77\%$ for PadChest and $79\%$ for EMBED, across shift levels with 1,000 test images. The overall identification accuracy for mixed shifts across all shifts and datasets is $85\%$ (with 1000 test images).
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Refer to caption(e)
Refer to caption(f)
Refer to caption(g)
Refer to caption(h)
Refer to caption(i)
@@
Figure 5: Shift identification accuracy: prevalence shifts (a-d) and covariate shifts (e-i). Across all datasets, the shift identification framework is able to successfully detect and identify both prevalence shifts and covariate shifts with high accuracy. Identification accuracy is computed over 200 bootstrap samples.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Refer to caption(e)
@@
Figure 6: Shift identification accuracy: mixed covariate and prevalence shifts. (a-b) depicts mixed gender and prevalence shift, (c-e) show mixed acquisition and prevalence shift. Across all datasets, the shift identification framework is able to successfully detect and identify presence of both shifts with high accuracy. Identification accuracy is computed over 200 bootstrap samples.
Discussion
@@
Shift identification is critical for root cause analysis of AI performance drift. We propose a lightweight and practical shift identification framework, capable of detecting and identifying important real-world dataset shifts. Our results show that the proposed framework separates prevalence shifts from covariate shifts with high accuracy, as well as identifies cases of mixed shifts, across three different data modalities and various types of realistic shifts. By analysing common dataset shift detection paradigms, we find that different types of shifts require different types of shift detectors. Our analysis also demonstrates the importance of the choice of encoders for feature-based dataset shift detection. In particular, we find that encoders trained in a self-supervised manner yield features with substantially higher shift detection power than supervised counterparts. Maybe surprisingly, our results show that generic encoders trained in a self-supervised manner on natural images (ImageNet) provide highly discriminative features for medical image dataset shift detection, transportable across datasets. Following these findings, we evaluate a new dual dataset shift detector, combining shift detection signals from task model outputs tests and features from self-supervised encoders. This approach outperforms existing shift detectors. The consistency of shift detection performance across various types of shifts is crucial as the nature of the shift is, by definition, unknown at test time. We take this combined approach a step further to perform fine-grained shift identification, showing high accuracy in correctly identifying the type of shift present in the test set across all modalities, tasks and various levels of shift intensity. Our results demonstrate the practical value of the shift identification method, which does not require any training at test time, nor any ground truth labels or annotations on the test domain data. The use of a readily available self-supervised encoder trained on ImageNet data for feature extraction dispenses us from training any additional model for shift detection and identification purposes. Combining signals from both feature-based and model output-based shift detectors yields reliable and consistent detection and identification across all types of shifts. Importantly, our method not only separates covariate shifts from prevalence shifts but also reliably detects when both types of shifts are present in the test set, rendering the proposed method applicable to many real-world deployment scenarios.
@@
In practice, the proposed framework can be used as a continuous monitoring tool for any image-based, clinical AI model. Similar to other continuous monitoring tools [38, 4], the shift identification framework would run on the stream of incoming data, collecting test data on a rolling time window. The reference data should be a small dataset representative of the expected data distribution on which the model has been validated. Our results show that with at most 2000 reference images, a test set of only 500 images suffices to yield high shift detection and identification accuracy across all shifts and datasets ($>80\%$).
@@
In terms of limitations, we note that in the case of covariate shift, on its own, the proposed shift identification framework does not allow for a more fine-grained identification of the origin of shift, e.g. the distinction between population and acquisition shifts. To allow for an even more precise sub-type shift identification, integrating metadata statistics in the pipeline (e.g. as in multi-modal shift detection pipelines [38]) could complement the framework. Nevertheless, it is important to highlight that relying solely on metadata monitoring only enables the detection of shifts affecting the specific attributes collected at test time. For example, statistics on patient population, such as age distribution, gender distribution, may be recorded at deployment time and used to detect some sub-types of shifts. Should the metadata not be available at test time, one could use auxiliary models to predict attributes of interest from images directly (e.g. ethnicity[39]). However, solely relying on collected (or predicted) metadata may not capture all sources of covariate shifts and will not allow detection of prevalence shift. The proposed framework can help uncover shifts that are not detectable by means of simply tracking population metadata. Shift detection and identification can prompt further investigation and inform root cause analysis of AI performance degradation. In this context, our shift identification framework offers an important safeguard for deploying AI models in clinical practice.
Methods
Definitions of prevalence and covariate shift
Formally, let $X$ denote the input image and $Y$ denote the target (e.g. disease label). Label shift (or prevalence shift) occurs when $P_{ref}(Y)\neq P_{test}(Y)$, where $P_{ref}(Y)$ is the distribution of labels on the reference domain and $P_{test}(Y)$ on the test domain. Label shift [40] assumes that the conditional distributions as well the marginal input distributions are preserved, i.e. $P_{ref}(X|Y)=P_{test}(X|Y)$ and $P_{ref}(X)=P_{test}(X)$. Conversely, covariate shift occurs when $P_{ref}(X)\neq P_{test}(X)$, while the conditional distributions $P_{ref}(Y|X)$ and $P_{test}(Y|X)$ are preserved [40]. Acquisition and subpopulation shifts are both cases of covariate shifts as they directly affect image appearance, i.e. marginal data densities.
Dataset shift detection methods
Several paradigms have been proposed for dataset shift detection. The most widely used and simplest method to implement consists of comparing distributions of a classifier’s outputs between the reference and test domain, proposed by Rabanser et al. [19] and referred to as ‘Black Box Shift Detection’ (BBSD). Specifically, softmax model outputs (predicted probabilities) are collected for all samples in the reference and test sets. Then, for each class, a separate univariate Kolmogorov-Smirnov (K-S) test is run to determine if the class-wise predicted probabilities distributions differ between reference and test domain. A multiple-testing Bonferonni [41] correction is then applied to the individual tests to determine the overall significance of the test.
In the same study, Rabanser et al. [19] also propose another type of shift detector where reference and test data input distribution are compared using a feature-based approach. In this test, the input sample (image) first gets projected to a smaller dimension, e.g. through a pretrained neural network encoder and the shift is then measured using the ‘Maximum Mean Discrepancy’ permutation test originally proposed by [42]. The Maximum Mean Discrepancy measures the distance between two distributions $p$ and $q$ based on the distance between their mean embeddings. An unbiased estimate of the square of the MMD statistic can be computed via
$\widehat{MMD}=\frac{1}{m^{2}-m}\sum_{i=1}^{m}\sum_{j\neq i}^{m}\kappa(z_{i},z_{j})$
$+\frac{1}{n^{2}-n}\sum_{i=1}^{n}\sum_{j\neq i}^{n}\kappa(z^{\prime}_{i},z^{\prime}_{j})$
$-\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\kappa(z_{i},z^{\prime}_{j}),$
where $\{z_{i}\}_{i=1}^{m}\sim p$, and $\{z^{\prime}_{i}\}_{i=1}^{n}\sim q$ and $\kappa$ is a kernel over the embedding space. The p-value can then be obtained using a permutation test. In Rabanser et al. they proposed to use the RBF kernel $\kappa(z,\tilde{z})=e^{\frac{1}{\sigma}||z-\tilde{z}||^{2}}$, setting $\sigma$ as the median distance between all samples. An alternative way would be to explicitly learn the kernel [43] with ‘deep kernels’. The disadvantage of this approach is that the kernel needs to be learned for every single test set for which one wishes to run shift detection. In this work, we use the RBF kernel.
Additionally, we propose and evaluate a dual shift detector, combining BBSD and MMD outcomes (referred to as ‘Duo’). We first run each test independently, this yields $C$ p-values for the BBSD test (one per class), and one p-value for the MMD permutation test. To account for multiple testing, we then apply Bonferroni correction [41] on the $C+1$ p-values to get the overall dual test outcome.
Dataset shift identification
In this work, we propose a novel framework for identifying whether dataset shift is caused by prevalence shift, by covariate shift or by a mix of both. The process is summarised in Fig. 2. The approach can be divided into two stages. In the first step, we perform ‘standard’ dataset shift detection to separate the ‘shift’ from the ‘no shift’ cases. For this detection step, we use the proposed dual detection approach, combining signals from model outputs and self-supervised features to detect shifts (see Dataset shift detection methods). Provided that a shift has been detected in this first step, we add an additional ‘shift identification’ step to separate prevalence from covariate shifts. We start by estimating the label distribution in the test dataset using the density ratio estimator from state-of-the-art label adaptation method CPMCN [22] (see next section). Next, we under-sample the reference set to match the estimated test set prevalence shift. We then compare feature distributions between the prevalence-adjusted reference set and the test set. If adjusting the prevalence removes all significant differences between the reference and test set, the shift is attributed to prevalence shift. Otherwise, covariate shift is necessarily present. Finally, given that a covariate shift has been detected, we need to determine whether a prevalence shift is also present (mixed shift). To answer this question, we leverage the fact that model output-based and feature-based detectors exhibit different behaviours in function of the type of shifts. In particular, our analysis of shift detection methods showed that model output-based shift detection is best suited for prevalence shift detection. Hence, we compare model output distributions to separate cases of covariate shift only from cases of mixed shifts. Precisely, if output distributions were significantly different before adjusting the prevalence on the reference set, but differences disappear after adjusting the prevalence, we conclude that prevalence shift is also responsible for the observed shift, hence the observed shift is a mix of prevalence and covariate shift (mixed shift). Else, the shift is attributed to ‘covariate shift’ only.
Prevalence shift estimation and adaptation
Many studies have proposed algorithms to alleviate the detrimental effect of prevalence shift on machine learning models. Most methods follow the same over-arching idea: (i) estimate the density ratio $P_{ref}(Y)/P_{test}(Y)$; (ii) use this ratio to recalibrate the model [14, 23, 47, 22]. In this work, we are primarily interested in the first step, to estimate the prevalence in the test set in the absence of labels. Various methods have been proposed to estimate this density ratio, mostly differing by their computational costs [14, 23, 47, 22]. Here, we use the recently proposed state-of-the-art ‘class probability matching with calibration network’ (CPMCN) method by Wen et al. [22] to estimate this ratio. In this work, the authors estimate the density ratio $w\in\mathbb{R}^{C}$, with $w_{i}=\frac{P_{test}(Y=i)}{P_{ref}(Y=i)}$ by
$\hat{w}:=\operatorname*{arg\,min}_{w\in\mathbb{R}^{C}}\sum_{i=1}^{C}\left|\hat{P}_{ref}(Y=i)-\frac{1}{m}\sum_{x\in D_{t}}\frac{\hat{p}(i|x)}{\sum_{j}^{C}w_{j}\hat{p}(j|x)}\right|^{2}$
$\hat{P}_{test}(Y=i)=\hat{w}_{i}\cdot\hat{P}_{ref}(Y=i),\forall i\in C$
We follow this method to estimate the label distribution in the test set in our shift identification module.
Note that, once the density ratio has been estimated, it can be used to recalibrate model outputs to mitigate the effects of prevalence shifts on model performance. Precisely, re-calibrated probabilities can be computed by:
$\hat{q}(i|x)=\frac{\hat{w}_{i}\hat{p}(i|x)}{\sum_{j=1}^{m}\hat{w}_{j}\hat{p}(j|x)}$ (1)
where $\hat{p}(i|x)$ are the original model outputs and $\hat{q}(i|x)$ the recalibrated output, for sample $x$ and class $i$.
Self-supervised learning
Self-supervised learning aims to learn meaningful representations of images without relying on labelled data [21, 48, 49]. Two prominent paradigms for learning such representations are contrastive learning and predictive coding methods. In contrastive learning, models are trained to learn similar representations for semantically related pairs of image views (positive pairs) while pushing them away from the representations of unrelated inputs (negative pairs). A seminal work in this area is SimCLR [21], where different views are created using heavy random data augmentation pipelines and representations are optimised using a normalised temperature-scaled cross-entropy loss [21]. SimCLR has since been refined in numerous follow-up studies[48, 50, 51, 52], yet remains widely used. On the other hand, predictive coding approaches, such as Masked Auto Encoders [49], learn representations by masking parts of the input and training a neural network to reconstruct the missing regions, encouraging the network to build a fine-grained understanding of the semantic information present in the visible parts of the image.
Datasets
For chest radiography, we use two public datasets for our analysis: the RSNA Pneumonia dataset [31], a subset of the NIH Chest-Xray8 dataset [53] manually relabelled by expert radiologists for presence of pneumonia-like opacities. We use the original metadata from the NIH Chest-Xray8 dataset [53] to retrieve patient gender. We also use PadChest [30] a larger dataset containing multiple disease labels extracted from radiology reports. We here focus on the pneumonia label. Importantly, this dataset contains important metadata such as scanner information or the gender of the patient, allowing us to generate a wide range of shifts. For mammography, we use the EMBED dataset [32] a large mammogram dataset collected in the US, on 6 different scanners. Finally, for fundus imaging, we create ‘RETINA’ a multi-domain dataset by combining three different public datasets: the Kaggle Diabetic Retinopathy Detection [33] dataset, the Kaggle Aptos Blindness Detection dataset [34] and the Messidor-v2 dataset [35]. These datasets cover different regions of the world (India, France, US) but also with varying image acquisition devices: images from the Messidor-v2 are high-quality images, while many images in the Kaggle datasets are of lower quality (including phone pictures). Creating this multi-centre dataset allows us to simulate various domain shifts by varying the proportion of data from each ‘site’ (original dataset source). The task of interest for fundus images here is binary diabetic retinopathy (DR) classification for fundus images, where we classify images between referable DR (grades 2,3,4) and healthy/non-referable DR (grades 0,1).
Shift generation details
For prevalence shifts, we assume that the task of interest is pneumonia detection for chest x-ray images, breast density classification for mammography and binary diabetic retinopathy classification for fundus images. In terms of covariate shifts, we simulate different sub-types of shifts depending on the datasets at hand. For both chest radiography datasets, we first simulate sub-population shift in the form of gender shift where we vary the proportion of female patients in the test set. PadChest [30] additionally includes scanner labels, allowing the study of acquisition shift. The dataset features scans from two types of scanners, ’Phillips’ and ’Imaging’, with an initial 40%-60% distribution. To simulate different levels of acquisition shift, we vary the proportion of Phillips scans in the test set. Similarly, for the EMBED [32] mammography dataset, we simulate various levels of acquisition shift by varying the distribution of scanners in the test set. This dataset offers a complementary view to the chest x-ray datasets, addressing a multi-class problem (unlike the binary pneumonia detection task) and providing even more flexibility for simulating diverse levels of acquisition shifts (as it contains data from six different scanners). Note that in EMBED each standard mammography exam comprises 4 mammograms (left/right breasts and MLO/CC views). We excluded all exams that did not contain exactly four images and kept exactly one exam per patient. For this dataset, test set resampling was done at the exam level (as opposed to image-level sampling for other datasets). Finally, for our RETINA dataset, we simulate domain shifts by varying the proportion of samples coming from each underlying dataset (Aptos [34], Kaggle DR [33] and Messidor [35]).
Implementation details
For all datasets, we first define train, validation and test splits. The training set is reserved for training task models (pneumonia, density and DR classification). All task models are ResNet-50 models initialised with ImageNet weights. To evaluate the shift identification accuracy of the proposed frameworks, we follow standard evaluation practices from the dataset shift detection literature. Specifically, we repeat the following process 200 times: (i) sample a subset of size $N_{test}$ from the original test split according to the shift of interest, (ii) run the shift identification test, and (iii) record whether the shift is correctly identified. The final shift identification accuracy is computed as the proportion of times the shift is correctly identified out of the bootstrap samples. For each bootstrap repetition, we also re-sample the reference set (against which we measure the shift) by taking a random sample of $N_{ref}$ data points from the validation set (uniformly at random). For both chest radiography datasets, we used $N_{ref}=2,000$ images and $N_{test}\in\{100,250,500,1000\}$ images. For the RETINA dataset, we used $N_{ref}=1,000$ images and $N_{test}\in\{100,250,1000\}$ images. The case of EMBED is slightly different as we need to ensure that images belonging to the same exam are always sampled simultaneously. Hence, for this dataset sampling of all the reference sets, shifted test sets (and permutations in the MMD test [^1] Keeping the underlying structure of the data during permutation tests is important to avoid inflated type I error [54].) is done at the exam level and not at the image level. We assumed $N_{exams,ref}=1,000$ exams (i.e $4,000$ images) and $N_{exams,test}\in\{50,100,250\}$.
For each dataset, we use the task model for the BBSD test. For the MMD detection test, we first extract the embeddings from the last layer of the encoder, then project them onto the 32-first principle components and then use the RBF kernel (see Dataset shift detection methods). We compare various encoders for feature extraction, in particular encoders trained in a self-supervised manner. Self-supervised encoders were trained using the SimCLR [21] objective, with a ResNet-50 [55] architecture. Additionally, for the RETINA dataset, we compare with RetFound [37] a publicly available self-supervised foundation model trained with the Masked Auto Encoder objective on a large set of retinal images, based on a transformer architecture (ViT [56]). All the code necessary to reproduce our experiments is made publicly available. [^2] https://github.com/biomedia-mira/shift˙identification.
Acknowledgements
@@
M.R. is funded by an Imperial College London President’s PhD Scholarship and a Google PhD Fellowship. R.M. is funded through the European Union’s Horizon Europe research and innovation programme under grant agreement 101080302. C.J. is supported by Microsoft Research and the Engineering and Physical Sciences Research Council (EPSRC) through a Microsoft PhD Scholarship. B.G. acknowledges support from the Royal Academy of Engineering as part of his Kheiron Medical Technologies/RAEng Research Chair in Safe Deployment of Medical Imaging AI and the UKRI AI programme, and the EPSRC, for CHAI - Causality in Healthcare AI Hub [grant number EP/Y028856/1].
Competing interests
B.G. is a part-time employee of HeartFlow and Kheiron Medical Technologies. All other authors declare no competing interests.
Supplementary material
Appendix A Appendix A: Motivating example
In this section, our goal is to motivate the need for dataset shift identification, illustrating the usefulness of distinguishing between prevalence shift and covariate shift with a simple example on chest radiography data.
For this motivating example, we use data from the PadChest dataset [30] and focus on classifying images with pneumonia. We then generate various test sets with (i) prevalence shift, (ii) acquisition shift, (iii) both shifts and evaluate the model on these shifted datasets. Our goal is to illustrate why identifying the type of shift occurring is a crucial issue. Indeed, many domain adaptation methods are tailored for specific shifts; label shift adaptation, for example, aims at tackling the effect of prevalence shift only, often explicitly assuming that no other shift is occurring. In supplementary Fig. 1, we show the effect of dataset shift on model calibration before and after applying a post-hoc label shift adaptation algorithm in three settings: (i) only prevalence shift occurs, (ii) acquisition shift occurs instead, (iii) both shifts happen. For prevalence shift, we sampled a test set with 20% prevalence of pneumonia, shifted from the original 3% prevalence in PadChest (training and validation prevalence). For acquisition shift, we create a test set containing 90% of scans taken with Phillips scanner and 10% with Imaging scanner, a shift from the original 40%-60% scanner distribution in PadChest. The label shift adaptation algorithm used for this example is CPMCN [22] (see Prevalence shift estimation and adaptation). We can clearly see that all shifts cause some amount of model calibration deterioration. More importantly, in the prevalence shift case, CPMCN successfully restores model calibration on the shifted dataset, and the estimated prevalence matches the true test set prevalence. However, for the acquisition shift scenario, the effect of applying label shift correction is counter-productive: calibration is much worse after applying the re-calibration than before. In the case where both shifts are happening at the same time, the re-calibration also fails to restore the calibration. This is, of course, absolutely expected from the theory: label shift adaptation methods like the ones tested here were designed for the case where only prevalence shift is happening. However, this empirical observation confirms the main motivation of our study: it is crucial to be able to identify the cause of dataset shifts for safe deployment.
Figure 1: Why dataset shift identification matters: example of wrongly applying label shift adaptation algorithms in scenarios when covariate shift is present. Subplots from left to right show model calibration plots for three scenarios: (i) prevalence shift, (ii) acquisition shift, and (iii) both shifts for a pneumonia detection model on the PadChest dataset. The re-calibration is successful in the prevalence shift only settings but fails (or even harms) in the presence of covariate shift.
Appendix B False detection rate for different shift detectors
In this section, we measure the false detection rate of dataset shift in the absence of shift in the test set generating process. We verify that the false-positive rates of the output-based, feature-based and Duo detectors stay low across datasets and detectors, we find that false detection rates hover around the expected type-I error of the statistical tests 5% for most datasets and detection methods. The only exception is EMBED, where false positive rates are a little higher than expected for the output-based (and hence Duo) detector. This is due to the fact that - for this highly imbalanced classification problem - small randomly resampled test sets may lead to small differences in the observed cumulative distribution of predictions, even in the absence of a shift in the test set generation process.
Figure 2: False detection rate. Percentage of shift detected when resampling the test set without shift.

### v3

@@ 4.2 Shift identification performance
To perform shift identification, we leverage the fact that output-based and feature-based shift measures detect different types of shifts to precisely identify the shift present in the test dataset, following the decision logic detailed in Section 3.1 and illustrated in Fig. 2.
@@
In-depth evaluation results in Figs. 6 and 7 demonstrate that our novel shift identification framework is capable of distinguishing between prevalence shifts, covariate shifts and mixed shifts with high accuracy across all datasets and types of shifts. Overall, more subtle shifts are best detected with larger test sets whereas larger shifts can be detected with smaller test sets. For prevalence shifts, the average shift identification rate, across datasets and shift levels, is 85% with 500 test images and reaches 95% with 1000 test images (Fig. 6, top two rows).
@@
Figure 6: Shift identification accuracy: prevalence shifts (top two rows) and covariate shifts (bottom three rows). Across all datasets, the shift identification framework is able to successfully detect and identify both prevalence shifts and covariate shifts with high accuracy. Identification accuracy is computed over 200 bootstrap samples.
@@
Figure 7: Shift identification accuracy: mixed covariate and prevalence shifts. (Top row) depicts mixed gender and prevalence shift, (bottom two rows) show mixed acquisition and prevalence shift. Across all datasets, the shift identification framework is able to successfully detect and identify presence of both shifts with high accuracy. Identification accuracy is computed over 200 bootstrap samples.
Similarly, for covariate shifts caused by acquisition shifts, when using a test set size of 500 images (and 250 test exams on EMBED), the identification accuracy is greater than 80% for any shift level and dataset, with an average identification accuracy of 89% across datasets and shift levels (Fig. 6 bottom two rows). For the RETINA dataset, identification rates already reach 100% with a test set as small as 250 images. When the covariate shift is induced by gender shifts, covariate shift is detected with identification accuracies greater than 90% for both PadChest and RSNA Pneumonia, with a test set size of 500 images, for all but one shift level (Fig. 6). Results in Fig. 7, show that the framework is also able to accurately distinguish between cases of covariate shift only and cases of covariate and prevalence shifts. For these mixed shifts, shifts are identified with increasingly high accuracy as the test set size increases. With a test set of 1000 images, mixed gender and prevalence shifts are correctly identified as mixed shifts with an average accuracy of 97% for RSNA Pneumonia and 75% for PadChest. Mixed shifts induced by acquisition and prevalence shifts are detected with an average 100% accuracy for the RETINA dataset, 77% for PadChest and 79% for EMBED, across shift levels with 1000 test images. The overall identification accuracy for mixed shifts across all shifts and datasets is 85% (with 1000 test images).
5 Discussion
@@
By analysing common dataset shift detection paradigms, we find that different types of shifts require different types of shift detectors. Our analysis also demonstrates the importance of the choice of encoders for feature-based dataset shift detection. In particular, we find that encoders trained in a self-supervised manner yield features with substantially higher shift detection power than supervised counterparts. Maybe surprisingly, our results show that generic encoders trained in a self-supervised manner on natural images (ImageNet) provide highly discriminative features for medical image dataset shift detection, transportable across datasets. Following these findings, we evaluate a new dual dataset shift detector, combining shift detection signals from task model outputs tests and features from self-supervised encoders. This approach outperforms existing shift detectors. The consistency of shift detection performance across various types of shifts is crucial as the nature of the shift is, by definition, unknown at test time.
We take this combined approach a step further to perform fine-grained shift identification, showing high accuracy in correctly identifying the type of shift present in the test set across all modalities, tasks and various levels of shift intensity. Our results demonstrate the practical value of the shift identification method, which does not require any training at test time, nor any ground truth labels or annotations on the test domain data. The use of a readily available self-supervised encoder trained on ImageNet data for feature extraction dispenses us from training any additional model for shift detection and identification purposes. Combining signals from both feature-based and model output-based shift detectors yields reliable and consistent detection and identification across all types of shifts. Importantly, our method not only separates covariate shifts from prevalence shifts but also reliably detects when both types of shifts are present in the test set, rendering the proposed method applicable to many real-world deployment scenarios.
@@
In practice, the proposed framework can be used as a continuous monitoring tool for any image-based, clinical AI model. Similar to other continuous monitoring tools [28], the shift identification framework would run on the stream of incoming data, collecting test data on a rolling time window. The reference data should be a small dataset representative of the expected data distribution on which the model has been validated. Our results show that with at most 2000 reference images, a test set of only 500 images suffices to yield high shift detection and identification accuracy across all shifts and datasets ($>80\%$).
@@
In terms of limitations, we note that in the case of covariate shift, on its own, the proposed shift identification framework does not allow for a more fine-grained identification of the origin of shift, e.g. the distinction between population and acquisition shifts. To allow for an even more precise sub-type shift identification, integrating metadata statistics in the pipeline (e.g. as in multi-modal shift detection pipelines  [28]) could complement the framework. Nevertheless, it is important to highlight that relying solely on metadata monitoring only enables the detection of shifts affecting the specific attributes collected at test time. For example, statistics on patient population, such as age distribution, gender distribution, may be recorded at deployment time and used to detect some sub-types of shifts. Should the metadata not be available at test time, one could use auxiliary models to predict attributes of interest from images directly (e.g. ethnicity [14]). However, solely relying on collected (or predicted) metadata may not capture all sources of covariate shifts and will not allow detection of prevalence shift. The proposed framework can help uncover shifts that are not detectable by means of simply tracking population metadata. Shift detection and identification can prompt further investigation and inform root cause analysis of AI performance degradation.
Acknowledgments
@@
M.R. is funded by an Imperial College London President’s PhD Scholarship and a Google PhD Fellowship. R.M. is funded through the European Union’s Horizon Europe research and innovation programme under grant agreement 10108030. C.J. is supported by Microsoft Research and EPSRC through the Microsoft PhD Scholarship Programme. B.G. acknowledges support from the Royal Academy of Engineering as part of his Kheiron Medical Technologies/RAEng Research Chair in Safe Deployment of Medical Imaging AI.
Disclosure of interests
B.G. is part-time employee of DeepHealth. No other competing interests.